# Vivid-VR: Distilling Concepts from Text-to-Video Diffusion Transformer for Photorealistic Video Restoration

### v1

@@ 3.2 Concept Distillation Training Strategy
PSNR $\uparrow$ 20.84 20.18 19.68 22.33 22.51 18.49 20.47 21.05 20.18
SSIM $\uparrow$ 0.595 0.532 0.527 0.594 0.645 0.467 0.548 0.593 0.523
LPIPS $\downarrow$ 0.240 0.313 0.292 0.281 0.308 0.215 0.224 0.178 0.287
@@
PSNR $\uparrow$ 27.57 27.02 28.97 28.20 27.29 18.36 23.23 23.45 24.54
SSIM $\uparrow$ 0.857 0.816 0.873 0.826 0.855 0.643 0.757 0.777 0.761
LPIPS $\downarrow$ 0.187 0.208 0.158 0.196 0.167 0.259 0.167 0.142 0.243
@@
PSNR $\uparrow$ 23.02 21.57 23.24 22.31 22.92 20.11 22.17 23.49 21.31
SSIM $\uparrow$ 0.655 0.585 0.639 0.592 0.657 0.588 0.620 0.677 0.579
LPIPS $\downarrow$ 0.341 0.347 0.350 0.340 0.433 0.314 0.296 0.258 0.357
@@ 4.2 Quantitative Results
Table 1 presents quantitative comparisons on $6$ benchmark testsets. The proposed Vivid-VR significantly outperforms existing methods in no-reference metrics, achieving the best results in almost all metrics. At the same time, we also note its advantages in full-reference metrics appear less pronounced. We argue that this arises primarily from the inherent limitations of these metrics, which often fail to align with human perceptual preferences. For example, the LPIPS values of Figure 4(g1) and (i1) are $0.3566$ and $0.4297$, respectively, while Figure 4(i1) is more prefered by human. This phenomenon becomes particularly evident in generative restoration scenarios where severe input degradation allows multiple plausible high-quality outputs, making full-reference metrics inadequate for quality assessment. This has also been mentioned in [42] and has been noticed by quality assessment studies [2, 11, 10].

### v2

@@ 3.2 Concept Distillation Training Strategy
PSNR $\uparrow$ 23.19 21.86 21.02 23.01 24.18 24.80 24.08 26.07 21.73
SSIM $\uparrow$ 0.690 0.609 0.595 0.606 0.720 0.754 0.689 0.777 0.604
LPIPS $\downarrow$ 0.230 0.304 0.281 0.277 0.301 0.168 0.263 0.191 0.278
@@
PSNR $\uparrow$ 27.57 27.02 28.97 28.20 27.29 30.53 27.80 29.04 24.54
SSIM $\uparrow$ 0.857 0.816 0.873 0.826 0.855 0.894 0.848 0.884 0.761
LPIPS $\downarrow$ 0.187 0.208 0.158 0.196 0.167 0.101 0.148 0.117 0.243
@@
PSNR $\uparrow$ 23.02 21.57 23.24 22.31 22.92 24.10 22.46 24.00 21.31
SSIM $\uparrow$ 0.655 0.585 0.639 0.592 0.657 0.688 0.621 0.693 0.579
LPIPS $\downarrow$ 0.341 0.347 0.350 0.340 0.433 0.283 0.240 0.185 0.357
@@ 4.2 Quantitative Results
Table 1 presents quantitative comparisons on $6$ benchmark testsets. The proposed Vivid-VR significantly outperforms existing methods in no-reference metrics, achieving the best results in almost all metrics. At the same time, we also note its advantages in full-reference metrics appear less pronounced. We argue that this arises primarily from the inherent limitations of these metrics, which often fail to align with human perceptual preferences. For example, the LPIPS values of Figure 4(g1) and (i1) are $0.3112$ and $0.4297$, respectively, while Figure 4(i1) is more prefered by human. This phenomenon becomes particularly evident in generative restoration scenarios where severe input degradation allows multiple plausible high-quality outputs, making full-reference metrics inadequate for quality assessment. This has also been mentioned in [42] and has been noticed by quality assessment studies [2, 11, 10].